# A radial scalar product for Kerr quasinormal modes

### v3

@@ V.1 Direct integration
For $a=0.99$ (the central panel), the integration path appears to cross the branch cut many times. The top panel of Fig. 2 shows this in detail. There, the apparent crossing of branch cuts is an artifact of presentation: $\xi(z)$ actually runs along all branches of the Riemann surface defined by $\text{W}(z)$, but in Fig. 1, only a single branch of $\text{W}(\xi)$ is shown Kuijlaars et al. (2003). This can be understood by noting that the transformed scaler-product, Eq. (35), uses an integration path that follows the real line, where there are no branch cuts. This is shown explicitly in the bottom panel of Fig. 2.
@@ V.2 Analytic continuation
[…]
@@
The Gamma functions is constructed to extend the factorial function to the complex numbers. The confluent hypergeometric function is a product of Gamma functions and a power series in z, and so it is straightforward to evaluate.
This parade of special functions has the following use: Every $\text{a}(\xi)$ and $\text{b}(\xi)$ of interest to Eq. (42) will be smooth, and so posses a series expansion,
@@
[…]
@@
In Eq. (47a), we have effectively defined $\xi^{p}\right>$ to be ${\langle{1}\,|\,{\xi^{p}}\rangle}$. In Eq. (47b), we have applied the definition of the scalar product, and in Eq. (47c), we have equated the integral over $\xi$ with the product of special functions found in Eq. (43), but with $\text{B}_{0}\rightarrow\text{B}_{0}+p$. The result is that the scalar product of any two smooth functions may be equated with a sum over confluent hypergeometric terms,
[…]
×$\hskip 56.9055pt\times\,U(1+\text{B}_{0}+j+k,-\text{B}_{1},-\text{B}_{2})\;.$
In practice, it is useful to precompute monomial moments using Eq. (V.2), and then compute scalar products using Eq. (46). The monomial moments are defined up to an overall constant. Henceforth, we chose to normalize all $\xi^{p}\right>$ such that $\xi^{1}\right>=1$,
[…]
@@ VI Limitations of direct integration and analytic continuation
Direct integration, analytic continuation, and their application to monomial moments and confluent Heun polynomials are all subject to various caveats and limitations. Here, these facets are reviewed. At times, technical details are omitted for brevity, but particular attention is given to practical workarounds. While analytic continuation and direct integration are found to be complementary, analytic continuation is determined to be preferred.
@@ VI.1 Direct Integration: disallowed BH spins, and the zero frequency limit
×$\quad\quad\times\;\left(1-(\text{x}_{2}+1)z^{\text{x}_{0}}\right)^{-\text{B}_{0}-\text{B}_{1}-2}\;.$
@@
For physical applications, it happens that for any QNM $(\ell,m,n)$, having frequency
@@
In this, direct integration’s underlying limitation is one of mathematical perspective: it happens to be more robust to use analytic continuation via integration over closed contours, rather than singularity avoiding integration over finite paths. For example, Ref. Ronveaux and Arscott (1995) details how single and double Pochhammer contours may be used in similar mathematical settings. This approach makes use of standard theorems in complex analysis, and is ultimately equivalent to analytic continuation as outlined in Sec. V.2.
@@ VI.2 Analytic continuation and limitations of the Gamma function
It was found in Sec. V.2 that scalar products may be evaluated analytically using monomial moments, $\xi^{p}\right>$,
$\xi^{p}\right>\;$ $=\;\Gamma(1+\text{B}_{0}+p)\,U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})\;.$ (60a)
@@
$\xi^{p}\right>\;$ $\approx\;\frac{\Gamma(1+\text{B}_{0}+p)\,\Gamma(1+\text{B}_{1})}{\Gamma(2+\text{B}_{0}+p+\text{B}_{1})}\;,$ (60b)
@@
For either Eq. (60a) or Eq. (60b), $\xi^{p}\right>$ will be regular when instances of the Gamma function are regular. Since $\Gamma(z)$ is well known to have simple poles when $z$ is a negative integer, it follows that $\xi^{p}\right>$ can diverge when,
@@
In Eq. (VI.2), $\{z\in\mathbb{Z}\,|\,z<0\}$ denotes the set of all negative integers. The practical implications of Eq. (VI.2) may be investigated by rewriting $\text{B}_{0}$ and $\text{B}_{1}$ in terms of physical parameters,
@@
In Eq. (VI.2), $z_{0}$ and $z_{1}$ are natural numbers, and we have used Eq. (8), Eqs. (III.4) and (IV.1) to expand in terms of basic physical quantities. Equation (62a) is equivalent to Eq. (61a), and Eq. (62b) is equivalent to Eq. (61b).
@@
Figure 3: Example distributions of absolute monomial moments, $|\langle\xi^{k}\rangle|$ v.s. monomial order, $k$, for the $(\ell,m)=(2,2)$ QNM at BH dimensionless spin of $a/M=0.7$. Left, central and right panels show results for $n$ of 00, $2$ and $6$, respectively. The corresponding QNM frequencies are $M\tilde{\omega}=0.5326-0.0808i$, $M\tilde{\omega}=0.4999-0.4123i$, and $M\tilde{\omega}=0.4239-1.0954i$ respectively. At focus is the effect of negating the spin weight, $s$, for increasing overtone label. For all cases shown, $|\langle\xi\rangle|=1$ has been enforced. Closed circles show points for $s=+2$, and open circles show points for $s=-2$.
@@ VII Pedagogical Applications
[…]
@@
In this section, we will first apply the scalar product to the monomial moments for different QNM values of the frequency $\tilde{\omega}$, paying particular attention to the effect of spin weight, $s$. For a broad range of scenarios, monomial moments with $s=2$ decrease monotonically with monomial order, while this is less so the case for $s=-2$. We will then apply the scalar product to the confluent Heun polynomials.
@@
Here, we take a different perspective: We will a priori define $\tilde{\omega}{}$ to be any QNM frequency  [^13] In the language of Ref. Ronveaux and Arscott (1995), we choose to focus on Class I of confluent Heun polynomials.. This requires that, in effect, we redefine the potential of $\mathcal{L}_{\xi}$ such that its eigenfunctions are polynomials. The consequences of this choice will have a significant impact on forthcoming results and discussion.
@@ VII.1 Monomial moments
For many common scalar products, monomial moments $\xi^{p}\right>$ decrease monotonically with $p$. This is easily demonstrated by considering a hypothetical constant weight function:
@@
[…]
$\xi^{p}\right>_{\text{hypothetical}}\;$ $=\;(p+1)^{-1}\sim p^{-1}\;.$ (65c)
@@
$\xi^{p}\right>\;$ $=\;\Gamma(1+\text{B}_{0}+p)\,U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})\;$ (66)
The confluent hypergeometric function, $U$, is defined as a series expansion in its third input DLMF (a). Thus some understanding of $\xi^{p}\right>$ may be found by examining $U$ expanded in powers of $\text{B}_{2}$. Examination of the zeroth order term happens to be sufficient for qualitative understanding of how $\xi^{p}\right>$ behaves when $p$ is large. For precise and predictive understanding, one would have to extend the forthcoming analysis to higher orders. This will not be done here so that the following discussion remains fairly simple.
@@
$\xi^{p}\right>\;$ $\approx\;\frac{\Gamma(1+\text{B}_{0}+p)\,\Gamma(1+\text{B}_{1})}{\Gamma(2+\text{B}_{0}+p+\text{B}_{1})}$ (68a)
@@
To go from Eq. (68a) to Eq. (68b), we have used the fact that, as $p\rightarrow\infty$, $\Gamma(p+\alpha)\,\sim\,p^{\alpha}\,\Gamma(p)$ Abramowitz and Stegun (1964). The physics of the situation is revealed by using Eqs. (IV.1) and (III.4) to rewrite $\text{B}_{1}$ in terms of $M$, $\tilde{\omega}$ and $s$,
$\xi^{p}\right>\;$ $\sim\;p^{-1-2s+4iM\tilde{\omega}}\;$ (69a)
@@
In Eq. (69a), we see that the previous result of $p^{-1}$ is effectively generalized to depend on the spin weight, $s$, and the dimensionless frequency, $M\tilde{\omega}$. Since the net exponent of $p$ is now complex valued, it is the real part of the exponent that determines the large-$p$ behavior of |$\xi^{p}\right>|$ |. In Eq. (69b), this is shown explicitly by using $\tilde{\omega}=w-i/\tau$. Recall that $\tau$ is strictly positive as is needed for time domain stability.
The presence of $e^{4iM\omega\,\ln(p)}$ in Eq. (69b) means that, in general, the real and imaginary parts of $\xi^{p}\right>$ will trace out a logarithmic spiral.
@@
    $\xi^{p}\right>|_{s=+2}\;\sim\;\ p^{-8}\,\left<\xi^{p}\right>|_{s=-2}$ (70)
@@
    Lastly, the presence of $4M/|\tau|$ in Eq. (69b) means that, as $\tau$ decreases, the rate of change of $\xi^{p}\right>$ will typically increase. This is perhaps most relevant for the QNM overtones, for which $|\tau|$ decreases as the overtone label, $n$, increases Zimmerman and Mark (2016); Berti et al. (2003).
@@
    In the left panel ($n=0$), and for $s=-2$, we see that the monomial moments |$\xi^{p}\right>|$ | decrease monotonically with $p$. However, for $s=-2$ and $n\in\{2,6\}$, moments are non-monotonic, and at $n=6$, they increase as $p$ increases.
@@
    In each panel, we see that monomial moments for spin weight $s=+2$ are significantly depressed relative to those for $s=-2$.
@@
The fact that absolute monomial moments, |$\xi^{p}\right>|$ |, are not strictly decreasing with $p$ has implications for the numerical behavior of scalar products. Naively, it might seem that, if |$\xi^{p}\right>|$ | increases with $p$, then scalar products might diverge. This happens to not be the case because the monomial moments are generally complex valued.
@@
Figure 4: The first four confluent Heun polynomials for $p=k$, spin weight $s=-2$, black hole dimensionless spin $a/M=0.86$, and $\tilde{\omega}$ defined by the $(\ell,m,n)=(3,3,1)$ QNM. The corresponding QNM frequency is $M\tilde{\omega}=0.9840-0.2140i$. Polynomial orders $p\in\{0,1,2,3\}$ are shown as a function of the physically valued fractional radius $\xi$. The top left and right panels shows the real and imaginary parts of each polynomial, respectively. Bottom left and right panels show the corresponding polynomial amplitudes and phases, respectively.
We have previously noted that the real and imaginary parts of $\xi^{p}\right>$ will trace out a logarithmic spiral in $p$. This implies that linear combinations of monomial moments will, in general, have a telescoping quality: as $p$ increases, successive terms will interfere destructively, meaning that although |$\xi^{p}\right>|$ | may increase, its net contribution must be considered along with adjacent terms. This telescoping quality behoves the use of arbitrary precision (e.g. mpmath mpmath development team (2023)) when evaluating moments by analytic continuation [^14] Leaver came to the same conclusion during his analysis of series expansions in $\xi$, which he refers to as Jaffé-type expansion Leaver (1986). Leaver (1986); Ronveaux and Arscott (1995). However, this quality also prompts a question.
@@ VII.2 Confluent Heun polynomials
Since the operator given in Eq. (VII.2) is of the confluent Heun type, its eigen-solutions are the aptly named confluent Heun functions Ronveaux and Arscott (1995); Magnus et al. (2021).
@@
    Classical polynomials are typically defined to be eigenfunctions of a purely differential operator (i.e. one with no $\partial_{\xi}^{0}$ term). For the confluent Heun operator, Eq. (VII.2), this purely differential part is denoted by Eqs. (71b) and (71c), and repeated below in Eq. (72),

### v4

@@ V.1 Direct integration
For $a=0.99$ (the central panel), the integration path appears to cross the branch cut [^13] For complex numbers $y$, We define the principle branch of $\log(y)$ by taking $\arg(y)\in(-\pi,\pi)$ and place the branch cut on the line defined by $\Re(y)<0$. As seen in Fig. 1, this convention results in a branch cut along $Re\xi<0$ (for powers of $\xi$), and another along $Re\xi>1$ (for powers of $1-\xi$). many times. The top panel of Fig. 2 shows this in detail. There, the apparent crossing of branch cuts is an artifact of presentation: $\xi(z)$ actually runs along all branches of the Riemann surface defined by $\text{W}(z)$, but in Fig. 1, only a single branch of $\text{W}(\xi)$ is shown Kuijlaars et al. (2003). This can be understood by noting that the transformed scaler-product, Eq. (35), uses an integration path that follows the real line, where there are no branch cuts. This is shown explicitly in the bottom panel of Fig. 2.
@@ V.2 Analytic continuation
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;=\;\int_{0}^{1}\,\text{a}(\xi)\,\text{b}(\xi)\;\;\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi\;.$ (42)
@@
The Gamma function is constructed to extend the factorial function to the complex numbers. The confluent hypergeometric function is a product of Gamma functions and a power series in z, and so it is straightforward to evaluate.
This parade of special functions has the following use: Every $\text{a}(\xi)$ and $\text{b}(\xi)$ of interest to Eq. (42) will be smooth, and so possesses a series expansion,
@@
$\left<\xi^{p}\right$>\; $=\;\int_{0}^{1}\,\xi^{p}\,\text{W}(\xi)\;d\xi$ (47a)
$=\;\int_{0}^{1}\,\xi^{\text{B}_{0}+p}\,(1-\xi)^{\text{B}_{1}}\,e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi$ (47b)
@@
In Eq. (47a), we have effectively defined $\left<\xi^{p}\right$> to be $\langle{1}\,|\,{\xi^{p}}\rangle$}. In Eq. (47b), we have applied the definition of the scalar product, and in Eq. (47c), we have equated the integral over $\xi$ with the product of special functions found in Eq. (43), but with $\text{B}_{0}\rightarrow\text{B}_{0}+p$. The result is that the scalar product of any two smooth functions may be equated with a sum over confluent hypergeometric terms,
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;$ $=\;e^{\text{B}_{2}}\,\sum_{j,k}\;\text{a}_{j}\,\text{b}_{k}\,\Gamma(\text{B}_{0}+j+k+1)\,$ (48)
$\hskip 56.9055pt\times\,U(1+\text{B}_{0}+j+k,-\text{B}_{1},-\text{B}_{2})\;.$
In practice, it is useful to precompute monomial moments using Eq. (47), and then compute scalar products using Eq. (46). The monomial moments are defined up to an overall constant. Henceforth, we chose to normalize all $\left<\xi^{p}\right$> such that $\left<\xi^{1}\right>=1$,
$\left<\xi^{p}\right>\;\leftarrow\;\left<\xi^{p}\right>/\left<\xi^{1}\right$>\;. (49)
@@ VI Limitations of direct integration and analytic continuation
Direct integration, analytic continuation, and their application to monomial moments and confluent Heun polynomials are all subject to various caveats and limitations. Here, these facets are reviewed. At times, technical details are omitted for brevity, but particular attention is given to practical workarounds. While analytic continuation and direct integration are equivalent when their respective caveats are satisfied, they are otherwise found to be complementary. Ultimately, analytic continuation is determined to be easier to use in practice.
@@ VI.1 Direct Integration: disallowed BH spins, and the zero frequency limit
$\quad\quad\times\;\left(1-(\text{x}_{2}+1)z^{\text{x}_{0}}\right)^{-\text{B}_{0}-\text{B}_{1}-2}\;.$
@@
For physical applications, it happens that for any QNM with $(\ell,m,n)$, having frequency
@@
In this, direct integration’s underlying limitation is one of mathematical perspective: it happens to be more robust to use analytic continuation via integration over closed contours, rather than singularity avoiding integration over finite paths as has been discussed above. For example, Ref. Ronveaux and Arscott (1995) details how single and double Pochhammer contours may be used in similar mathematical settings. This approach makes use of standard theorems in complex analysis, and is ultimately equivalent to analytic continuation as outlined in Sec. V.2.
@@ VI.2 Analytic continuation and limitations of the Gamma function
It was found in Sec. V.2 that scalar products may be evaluated analytically using monomial moments, $\left<\xi^{p}\right$>,
$\left<\xi^{p}\right$>\; $=\;\Gamma(1+\text{B}_{0}+p)\,U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})\;.$ (60a)
@@
$\left<\xi^{p}\right$>\; $\approx\;\frac{\Gamma(1+\text{B}_{0}+p)\,\Gamma(1+\text{B}_{1})}{\Gamma(2+\text{B}_{0}+p+\text{B}_{1})}\;,$ (60b)
@@
For either Eq. (60a) or Eq. (60b), $\left<\xi^{p}\right$> will be regular when instances of the Gamma function are regular. Since $\Gamma(z)$ is well known to have simple poles when $z$ is a negative integer, it follows that $\left<\xi^{p}\right$> can diverge when,
@@
In Eq. (61), $\{z\in\mathbb{Z}\,|\,z<0\}$ denotes the set of all negative integers. The practical implications of Eq. (61) may be investigated by rewriting $\text{B}_{0}$ and $\text{B}_{1}$ in terms of physical parameters,
@@
In Eq. (62), $z_{0}$ and $z_{1}$ are natural numbers, and we have used Eq. (8), Eqs. (16) and (24) to expand in terms of basic physical quantities. Equation (62a) is equivalent to Eq. (61a), and Eq. (62b) is equivalent to Eq. (61b).
@@
Figure 3: Example distributions of absolute monomial moments, $|\langle\xi^{k}\rangle|$ v.s. monomial order, $k$, for the $(\ell,m)=(2,2)$ QNM at BH dimensionless spin of $a/M=0.7$. Left, central and right panels show results for $n$ of 0, 22 and $6$, respectively. The corresponding QNM frequencies are $M\tilde{\omega}=0.5326-0.0808i$, $M\tilde{\omega}=0.4999-0.4123i$, and $M\tilde{\omega}=0.4239-1.0954i$ respectively. At focus is the effect of negating the spin weight, $s$, for increasing overtone label. For all cases shown, $|\langle\xi\rangle|=1$ has been enforced. Closed circles show points for $s=+2$, and open circles show points for s=2s=-2.
@@ VII Pedagogical Applications
${\langle{\text{a}}\,|\,{\text{b}}\rangle}\;=\;\int_{0}^{1}\,\text{a}(\xi)\,\text{b}(\xi)\;\;\xi^{\text{B}_{0}}(1-\xi)^{\text{B}_{1}}e^{\frac{\text{B}_{2}}{1-\xi}}\;d\xi\;.$ (63)
@@
In this section, we will first apply the scalar product to the monomial moments for different QNM values of the frequency $\tilde{\omega}$, paying particular attention to the effect of spin weight, $s$. For a broad range of scenarios, monomial moments with $s=2$ decrease monotonically with monomial order, while this is less so the case for s=2s=-2. We will then apply the scalar product to the confluent Heun polynomials.
@@
Here, we take a different perspective: We will a priori define $\tilde{\omega}{}$ to be any QNM frequency  [^15] In the language of Ref. Ronveaux and Arscott (1995), we choose to focus on Class I of confluent Heun polynomials.. This requires that, in effect, we redefine the potential of $\mathcal{L}_{\xi}$ such that its eigenfunctions are polynomials. The consequences of this choice enable confluent Heun polynomials to be defined for arbitrary frequencies; in turn, this will have a significant impact on forthcoming results and discussion.
@@ VII.1 Monomial moments
For many common scalar products, monomial moments $\left<\xi^{p}\right$> decrease monotonically with $p$. This is easily demonstrated by considering a hypothetical constant weight function:
@@
${\langle{\text{a}}\,|\,{\text{b}}\rangle}_{\text{hypothetical}}\;$ $=\;\int_{0}^{1}\,\text{a}(\xi)\,\text{b}(\xi)d\xi\;,$ (65b)
$\left<\xi^{p}\right>_{\text{hypothetical}}\;$ $=\;(p+1)^{-1}\sim p^{-1}\;.$ (65c)
@@
$\left<\xi^{p}\right$>\; $=\;\Gamma(1+\text{B}_{0}+p)\,U(1+\text{B}_{0}+p,-\text{B}_{1},-\text{B}_{2})\;$ (66)
The confluent hypergeometric function, $U$, is defined as a series expansion in its third input DLMF (a). Thus some understanding of $\left<\xi^{p}\right$> may be found by examining $U$ expanded in powers of $\text{B}_{2}$. Examination of the zeroth order term happens to be sufficient for qualitative understanding of how $\left<\xi^{p}\right$> behaves when $p$ is large. For precise and predictive understanding, one would have to extend the forthcoming analysis to higher orders. This will not be done here so that the following discussion remains fairly simple.
@@
$\left<\xi^{p}\right$>\; $\approx\;\frac{\Gamma(1+\text{B}_{0}+p)\,\Gamma(1+\text{B}_{1})}{\Gamma(2+\text{B}_{0}+p+\text{B}_{1})}$ (68a)
@@
To go from Eq. (68a) to Eq. (68b), we have used the fact that, as $p\rightarrow\infty$, $\Gamma(p+\alpha)\,\sim\,p^{\alpha}\,\Gamma(p)$ Abramowitz and Stegun (1964). The physics of the situation is revealed by using Eqs. (24) and (16) to rewrite $\text{B}_{1}$ in terms of $M$, $\tilde{\omega}$ and $s$,
$\left<\xi^{p}\right$>\; $\sim\;p^{-1-2s+4iM\tilde{\omega}}\;$ (69a)
@@
[…]
The presence of $e^{4iM\omega\,\ln(p)}$ in Eq. (69b) means that, in general, the real and imaginary parts of $\left<\xi^{p}\right$> will trace out a logarithmic spiral.
@@
    $\left<\xi^{p}\right>|_{s=+2}\;\sim\;\ p^{-8}\,\left<\xi^{p}\right>|_{s=-2}$ (70)
@@
    Lastly, the presence of $4M/|\tau|$ in Eq. (69b) means that, as $\tau$ decreases, the rate of change of $\left<\xi^{p}\right$> will typically increase. This is perhaps most relevant for the QNM overtones, for which $|\tau|$ decreases as the overtone label, $n$, increases Zimmerman and Mark (2016); Berti et al. (2003).
@@
[…]
@@
    In each panel, we see that monomial moments for spin weight $s=+2$ are significantly depressed relative to those for s=2s=-2.
@@
[…]
@@
Figure 4: The first four confluent Heun polynomials for $p=k$, spin weight s=2s=-2, black hole dimensionless spin $a/M=0.86$, and $\tilde{\omega}$ defined by the $(\ell,m,n)=(3,3,1)$ QNM. The corresponding QNM frequency is $M\tilde{\omega}=0.9840-0.2140i$. Polynomial orders $p\in\{0,1,2,3\}$ are shown as a function of the physically valued fractional radius $\xi$. The top left and right panels shows the real and imaginary parts of each polynomial, respectively. Bottom left and right panels show the corresponding polynomial amplitudes and phases, respectively.
[…]
@@ VII.2 Confluent Heun polynomials
Since the operator given in Eq. (71) is of the confluent Heun type, its eigen-solutions are the aptly named confluent Heun functions Ronveaux and Arscott (1995); Magnus et al. (2021).
@@
    Classical polynomials are typically defined to be eigenfunctions of a purely differential operator (i.e. one with no $\partial_{\xi}^{0}$ term). For the confluent Heun operator, Eq. (71), this purely differential part is denoted by Eqs. (71b) and (71c), and repeated below in Eq. (72),